Calculate the pH (give your answers to 2.d.p) of the following solution dilutions.

(a) 50 cm3 of 1.00 mol dm–3 hydrochloric acid is added to 200 cm3 of water.

(b) 100 cm3 of 1.00 mol dm–3 Ba(OH)2 is added to 400 cm3 of water at 298 K


1
Expert's answer
2020-11-26T05:09:09-0500

a)     C1V1=C2V2

C1=1.00moldm-3

V1= 50 cm3

C2=?

V2= (50+200) cm3

    =250 cm3

1 ×50= 250 C2

C2=50/250

   =0.2M

[H+] = 0.2M

pH=-log (0.2)

    =0.6990

    =0.70 (2dp)


b)     C1=1. 00moldm-3

V1=100 cm3

C2=?

V2= (100+400) cm3

    =500 cm3

100×1=500 C2

C2= 100/500

C2=0.2MBa(OH)2

Ba(OH)2(aq) =Ba2+(aq) +2OH-(aq)

[OH-]= 2×0.2= 0.4M

pOH=-log[OH-]

       =-log [0.4]

       =0.3979

pH=14-0.3979

     =13.6021

     =13.60 (2dp)